[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Aug 04, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Aug 4 20:17:17 UTC 2016


HSU Shift Log for Thu Aug 04 2016 16:17:10 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS/AWIPS2 issues:

Other Techincal issues: Internal radar loops from Mexico are not updating  
on snook. It's possible that the URLs where the images are stored changed  
and we need to update the script that pulls those in.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160804/a6e22588/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list